# SDTagNet: Leveraging Text-Annotated Navigation Maps for Online HD Map Construction

### v1

@@ 3 Method
An overview of SDTagNet is shown in Figure 1. Beginning with an introduction of annotations in SD maps (Section 3.1), we describe the two main components of SDTagNet: The NLP tag embedding module (Section 3.2) and the SD map encoder (Section 3.3).
@@ 3.1 Annotations in SD Maps
Second, SD maps are often considered sparse in content, leading to previous approaches only considering roads, e.g. by filtering for the highway tag in OpenStreetMap (OSM). As illustrated in Figure 2, existing methods neglect that SD maps, such as those made widely available by the OSM project [21], contain many more kinds of elements. They contain points like traffic lights, bus stops, traffic signs, etc. as well as many ways beyond roads, like pedestrian paths, building borders, etc. Virtual elements, so-called relations, even denote semantic or topological links between two elements, e.g. that no left turns are allowed.
Finally, there is a wide variety of descriptive text information annotated to many of those elements. This valuable knowledge has not yet been used to enhance online HD map construction. For instance, OSM maps contain so-called tags, i.e. key-value pairs, such as name: Park Avenue, oneway: yes, or lanes: 2. The last two examples already show that the textual information is highly informative for the task of online HD map construction. In total, at the point of writing, the global OSM map contains around 100k different keys and 168M values. This vast amount of tags as well as their unstructured nature obviously cannot be exploited by handcrafted methods. Instead, we propose an NLP encoder to optimally interpret the full information contained in SD maps for the goal of HD map construction.
@@ 3.2 NLP Tag Embeddings
Since a very large dataset for the use case exists as well with OpenStreetMap [21], we decided to train our embedding model from scratch. This also affords us more control of model hyperparameters such as embedding size and is necessary for reaching real-time performance.
For our pre-training we adapt established methods and design a self-supervised contrastive learning objective based on the multiple negatives ranking loss [7]. Our customized objective is motivated by the fact that many annotation artifacts and semantically irrelevant tags exist in crowdsourced maps. For instance, many elements in OSM contain IDs and other cataloging information of national geodetic reference systems that are not informative for a downstream task. Our customized loss forces the model to ignore these irrelevant tags, which we identified using the main OSM tag database.
Figure 3 shows an example of a training batch for the customized loss. Using the dataset of all unique tagsets from the OpenStreetMap planet map, the goal of the model is to minimize the distance of two tagsets with the same semantically meaningful tags while maximizing the distance towards other tagsets. A large batch size of negative samples is needed to keep the training stable [7], 5120 in our case. We randomly sample 20 pairs of positive samples per unique relevant tagset for the training data and choose a training time of 4 epochs to ensure convergence. To facilitate further research in this area, the pre-trained encoder together with the extracted OSM annotation dataset used for training will be released alongside the main model.
@@ 3.3 SD Map Encoder
Following the NLP tag embedding, the SD map encoder combines these embeddings with the associated points and the scene-level context through a transformer module. The SD map encoder module is based on the architecture proposed by SMERF [20], however with key changes that significantly improve performance in the online HD map construction task and allow for much more flexibility in its input representation. A detailed visualization of the SD map encoder design with all its components can be found in Figure 4. All changes are investigated individually in LABEL:subsubsec:ablation_studies and we show that their combined application significantly increases performance.
@@
In SMERF [20], one query token represents one polyline, which is misaligned with the detection transformer architectures in HD map construction using one detection query for each point instead [14, 15, 3, 4, 29]. To align the token design with the transformer architecture, we choose one query token per point having also the benefit of an easier introduction of additional element types of points and relations. As all polyline elements in vectorized HD map construction architectures are detected with a fixed point number, we also resample all SD map polylines to the fixed point number of 10. Furthermore, we adopt the sin/cos positional encoding for point coordinates from SMERF [20] in our encoder and concatenate the tag embedding of the respective map element to it.
@@
While point-level queries better align the element representations and provide additional flexibility, they introduce a problem: Without additional changes, context for the model is lost as to which element an individual point belongs to, for example, in the case of polyline features. We resolve this issue by taking inspiration from the field of graph transformers, where [10] show orthogonal random features (ORF) to be strong identifier features in the context of graph nodes and edges.
Following [10] and [28], we retrieve ORF features from the rows of the random orthogonal matrix $\mathbf{Q}\in\mathbb{R}^{n\times n}$, created from the QR decomposition of a random Gaussian matrix $\mathbf{G}\in\mathbb{R}^{n\times n}$. Each element is then assigned one ORF feature as its identifier, also visualized in Figure 4. As in [10], graph nodes, or in our case, point features and polyline features, have the same ORF feature vector stacked twice in their query. Relations or graph edges are then represented with the two ORF features of the respective graph nodes or SD map elements. This simple graph representation method for the SD map encoder is therefore not only a identifier for map element context, but also vastly broadens the element types that can be used for prior information.
@@ 4 Experiments
We evaluate our method on the Argoverse 2 [24] and nuScenes [1] datasets, viewing Argoverse 2 as the main dataset for evaluation. We discuss our selected datasets and metric in Section 4.1 and the implementation and baselines in Section 4.2. Section 4.3 contains the performance results of SDTagNet in comparison with existing methods. Ablation studies to assess the contributions of individual components in the SD map encoder and the optimization strategies applied to the NLP tag embedding module can be found in Section A.1.
@@ 4.1 Datasets and Metric
[…]
@@ 4.2 Implementation Details and Baselines
Our baseline SD map encoding approaches are the two recent works PMapNet [9] and SMERF [20] that represent the two general existing approaches towards SD map encoding: Rasterization into a BEV image in the case of PMapNet [9] and encoding of the vectorized polylines via a transformer module in the case of SMERF [20]. To minimize the influence from other priors like temporal information, we select the popular single-shot architecture MapTRv2 [15] as base model. For fair comparison, all SD map encoder modules have been ported to the MapTRv2 base architecture, thus isolating any performance differences to the SD map encoder module. All models are trained on 4 H100 Nvidia GPUs for 24 epochs on Argoverse 2 and 110 epochs for nuScenes.
Both baselines are evaluated in two settings: With their original input data settings and with all information that is available in the OSM map, to the extent that the encoding method permits its integration. For PMapNet [9], we add the more detailed road label classes from SMERF to the raster image and for SMERF [20] we introduce node features that are not used in the original work. This additionally eliminates any differences between available input data and compares SDTagNet with the fairest versions of existing SD map encoding methods. A detailed discussion of baseline implementation details and hyperparameters can be found in Section A.2.
@@ 4.3 Online HD Map Construction Performance
In this section we first discuss our results on Argoverse 2 and nuScenes in the near range and far range setting in comparison with previous work. Afterwards, we present ablation studies that show the effectiveness of the proposed architectural changes and NLP tag embedding module design.
Results on Argoverse 2
Table 1 shows the results of SDTagNet on Argoverse 2 [24] in the near range and far range setting in comparison with the SD map encoding baselines PMapNet [9] and SMERF [20], applied to the base architecture MapTRv2 [15].
The results show the strengths of the proposed SDTagNet architecture that fully exploits all information contained in SD prior maps, significantly outperforming existing methods even when beneficial changes are introduced to them. A qualitative comparison of SDTagNet with PMapNet and SMERF on Argoverse 2 in the far range setting is displayed in Figure 5.
@@
Dataset: Argoverse 2 Near Range (60 m $\times$ 30 m)
@@
Table 2 shows the evaluation results on the nuScenes dataset. The relative improvement of SDTagNet in the far range setting is even stronger, increasing performance compared to MapTRv2 by +4.1 mAP (+105%) and compared to existing methods by +2.1 mAP (+35%). This coincides with the general drop in mAP for all methods on the nuScenes geo split [17] and shows the benefits of our approach even on smaller datasets with less generalization capability for the main model. SMERF outperforms PMapNet here, indicating that these methods may be more susceptible to dataset differences than SDTagNet, which shows similar performance benefits in both cases. In the near range, the results mostly mirror those on Argoverse 2, with no method able to provide a significant performance gain. On nuScenes in particular, we observed some significant differences between the roads and structures contained in the SD map and the ground truth HD map (see Section 5). We hypothesize this is one of the main causes of the lacking performance in the near range, where the visibility of road features is much higher and the model is less reliant on SD map features.
@@
Dataset: nuScenes Near Range (60 m $\times$ 30 m)
@@ 5 Limitations
Even though SDTagNet shows strong performance on multiple datasets, we also observe a number of limitations. First, SDTagNet, like all SD map prior methods, assumes a congruency of SD map and HD map ground truth. Depending on the level of detail and quality of dataset ground truth and SD map, a number of map discrepancies can be observed, especially for a crowdsourced database like OSM. [19] and [9] have already shown discrepancies for Argoverse 2 and nuScenes respectively. We have also noted significant differences on nuScenes, particularly concerning structures and roads that are missing from the ground truth HD map. Figure 6(b) shows examples of a tunnel and smaller service roads that are not included in the nuScenes map, impacting model accuracy and training data quality.
Another observed limitation concerns the case of one-way roads. SDTagNet is able to correctly identify one-way roads from the SD map tag information, however in some cases the driving direction is predicted wrongly. An example of this is displayed in Figure 6(a). We hypothesize that, since OSM uses the polyline point order to encode one-way direction information, this is caused by the lack of polyline point order information in the element identifiers. Introduction of additional point order identifiers for polyline features could help remedy this limitation.
@@
Figure 6: Examples for observed limitations of SDTagNet in Argoverse 2 (left) and nuScenes (right). Figure 6(a): Even when correctly recognizing a one-way road, SDTagNet can fail to predict its correct driving direction. Figure 6(b): Discrepancies between dataset HD maps and OSM SD maps like missing tunnels (left) or small service roads (left, right).
@@ Appendix A Technical Appendices and Supplementary Material
The supplementary material focuses on three key aspects: Ablation studies on all individual components of SDTagNet (section A.1) and additional experiments on a non-geo split (section A.4), implementation details of the baselines and NLP pretraining (section A.2 and section A.3) as well as more qualitative examples and visualizations (section A.6).

### v2

@@ 3 Method
An overview of SDTagNet is shown in Figure˜1. Beginning with an introduction of annotations in SD maps (Section˜3.1), we describe the two main components of SDTagNet: The NLP tag embedding module (Section˜3.2) and the SD map encoder (Section˜3.3).
@@ 3.1 Annotations in SD Maps
Second, SD maps are often considered sparse in content, with previous approaches only considering roads, e.g. by filtering for the highway tag in OpenStreetMap (OSM). However, as shown in Figure˜2, SD maps like those widely available from the OSM [22] project contain many more kinds of elements. They contain points like traffic lights, bus stops, traffic signs, etc. as well as many ways beyond roads, like pedestrian paths, building borders, etc. Virtual elements, so-called relations, even denote semantic or topological links between two elements, e.g. that no left turns are allowed.
Finally, there is a wide variety of descriptive text information annotated to many of those elements. This valuable knowledge remains unexploited for online HD map construction. For instance, OSM maps contain so-called tags, i.e. key-value pairs, such as name: Park Avenue, oneway: yes, or lanes: 2. The last two examples already show that the textual information is highly informative for the task of online HD map construction. In total, at the point of writing, the global OSM map contains around 100k different keys and 168M values. This vast amount of tags as well as their unstructured nature obviously cannot be exploited by handcrafted methods. Instead, we propose an NLP encoder to optimally interpret the full information contained in SD maps for the goal of HD map construction.
@@ 3.2 NLP Tag Embeddings
Since a very large dataset for the use case exists as well with OpenStreetMap [22], we decided to train our embedding model from scratch. This also affords us more control of model hyperparameters such as embedding size and is necessary for reaching real-time performance. For our pre-training we adapt established methods and design a self-supervised contrastive learning objective based on the multiple negatives ranking loss [7]. Our customized objective is motivated by the fact that many annotation artifacts and semantically irrelevant tags exist in crowdsourced maps. For instance, many elements in OSM contain IDs and other cataloging information of national geodetic reference systems that are not informative for a downstream task. Our customized loss forces the model to ignore these irrelevant tags, which we identified using the main OSM tag database.
Figure˜3 shows an example of a training batch for the customized loss. Using the dataset of all unique tagsets from the OpenStreetMap planet map, the goal of the model is to minimize the distance of two tagsets with the same semantically meaningful tags while maximizing the distance towards other tagsets. A large batch size of negative samples is needed to keep the training stable [7], 5120 in our case. We randomly sample 20 pairs of positive samples per unique relevant tagset for the training data and choose a training time of 4 epochs to ensure convergence. To facilitate further research in this area, the pre-trained encoder together with the extracted OSM annotation dataset used for training will be released alongside the main model.
@@ 3.3 SD Map Encoder
Following the NLP tag embedding, the SD map encoder combines these embeddings with the associated points and the scene-level context through a transformer module. The SD map encoder module is based on the architecture proposed by SMERF [21], however with key changes that significantly improve performance in the online HD map construction task and allow for much more flexibility in its input representation. A detailed visualization of the SD map encoder design with all its components can be found in Figure˜4. All changes are investigated individually in Section˜4.3 and Section˜A.1 and we show that their combined application significantly increases performance.
@@
In SMERF [21], one query token represents one polyline, which is misaligned with the detection transformer architectures in HD map construction, that use one detection query for each point instead [15, 16, 3, 4, 30]. To align the token design, we choose one query token per point. This also has the benefit of an easier introduction of additional element types of points and relations. As all polyline elements in vectorized HD map construction architectures are detected with a fixed point number, we also resample all SD map polylines to the fixed point number of 10. Furthermore, we adopt the sin/cos positional encoding for point coordinates from SMERF [21] in our encoder and concatenate the tag embedding of the respective map element to it.
@@
While point-level queries better align the element representations and provide additional flexibility, they introduce a problem: Without additional changes, context for the model is lost as to which element a point belongs to, e.g. in the case of polyline features. We resolve this issue by taking inspiration from the field of graph transformers, where [10] show orthogonal random features (ORF) to be strong identifier features in the context of graph nodes and edges.
Following [10] and [29], we retrieve ORF features from the rows of the random orthogonal matrix $\mathbf{Q}\in\mathbb{R}^{n\times n}$, created from the QR decomposition of a random Gaussian matrix $\mathbf{G}\in\mathbb{R}^{n\times n}$. Each element is then assigned one ORF feature as its identifier, also visualized in Figure˜4.
As in [10], graph nodes, or in our case, point features and polyline features, have the same ORF feature vector stacked twice in their query. Points are represented as a single query of the SD map encoder with two identical ORF identifiers, a tag embedding, and positional encoding. Polylines are represented by a list of multiple point-like queries that all share the same two identical ORF identifiers. Relations are represented using pairs of ORF identifiers, each corresponding to one of the members of a relation edge, as done traditionally in graph-structured data with node IDs. This simple graph representation method for the SD map encoder is therefore not only a identifier for map element context, but also vastly broadens the element types that can be used for prior information.
@@ 4 Experiments
We evaluate our method on the Argoverse 2 [25] and nuScenes [1] datasets, viewing Argoverse 2 as the main dataset for evaluation. We discuss our selected datasets and metric in Section˜4.1 and the implementation and baselines in Section˜4.2. Section˜4.3 contains the performance results of SDTagNet in comparison with existing methods and ablation studies to assess the contributions of individual components in the SD map encoder. More detailed ablation studies regarding the optimization strategies applied to the NLP tag embedding module and fine-tuning training regimes can be found in Section˜A.1.
@@ 4.1 Datasets and Metric
Additionally, as SD maps, OpenStreetMap data corresponding to the frame-wise local HD map label is extracted from the planetary OSM database. We provide this data as SD map prior in two evaluation ranges for online HD map construction: near range ($60\text{\,}\mathrm{m}$ $30\text{\,}\mathrm{m}$) and far range ($120\text{\,}\mathrm{m}$ $60\text{\,}\mathrm{m}$), making sure to include all SD map elements that intersect that area. To the best of our knowledge, no previous method that evaluates on the increasingly important far range setting with a geographically disjoint train/val split [9, 30] combines these important evaluation settings. Comparing Table˜7 of Section˜A.4 with Table˜1 shows that SD priors aggravate the localization overfit observed by [18] when train and evaluation data overlaps geographically. We assume that this is due to the additional leaked information available to the network when using SD maps as prior. All evaluation is carried out with the standard mean Average Precision (mAP) metric with 0.5, 1.0 and 1.5 meters thresholds for the Chamfer distance. This matches the standard mAP metric used by our baseline architecture [16].
@@ 4.2 Implementation Details and Baselines
Our baseline SD map encoding approaches are the two recent works PMapNet [9] and SMERF [21] that represent the two general existing approaches towards SD map encoding: Rasterization into a BEV image in the case of PMapNet [9] and encoding of the vectorized polylines via a transformer module in the case of SMERF [21]. To minimize the influence from other priors like temporal information, we select the popular single-shot architecture MapTRv2 [16] as a base model. For fair comparison, all SD map encoder modules have been ported to the MapTRv2 base architecture, thus isolating any performance differences to the SD map encoder module. All models are trained on 4 H100 Nvidia GPUs for 24 epochs on Argoverse 2 and 110 epochs for nuScenes.
Both baselines are evaluated in two settings: With their original input data settings and with all information that is available in the OSM map, to the extent that the encoding method permits its integration. For PMapNet [9], we add the more detailed road label classes from SMERF to the raster image and for SMERF [21] we introduce node features that are not used in the original work. This additionally eliminates any differences between available input data and compares SDTagNet with the fairest versions of existing SD map encoding methods. A detailed discussion of baseline implementation details and hyperparameters can be found in Section˜A.2.
@@ 4.3 Online HD Map Construction Performance
In this section we first discuss our results on Argoverse 2 and nuScenes in the near range and far range setting in comparison with previous work. We also present ablation studies that show the effectiveness of the proposed architectural changes and NLP tag embedding module.
@@
Dataset: Argoverse 2 Near Range (60 m 30 m)
@@
Results on Argoverse 2
Table˜1 shows the results of SDTagNet on Argoverse 2 [25] in the near range and far range setting in comparison with the SD map encoding baselines PMapNet [9] and SMERF [21], applied to the base architecture MapTRv2 [16].
To compare SDTagNet with recent state-of-the-art non SD map prior methods, we also evaluated it in the more limited label setting of [3], the results of which are shown in Table˜2. SDTagNet maintains competitive performance in the near range, while significantly outperforming existing approaches in the far range, acheiving a +6.1+6.1 mAP gain vs. MapTracker [3]. This is despite SDTagNet being the only single-shot method in this comparison other than HIMap [33].
The results show the strengths of SDTagNet in fully exploiting all information contained in SD prior maps, significantly outperforming state-of-the-art non SD map prior methods and existing SD map encoders, even when beneficial changes are introduced to them. A qualitative comparison of SDTagNet with PMapNet and SMERF on Argoverse 2 in the far range setting is displayed in Figure˜5.
Ablation Studies
We conduct ablation studies on the Argoverse 2 dataset to assess the contributions of individual components in the SD map encoder, displayed in Table˜3. Only the full configuration of SDTagNet, combining all proposed modifications, yields significant performance gains compared to the SMERF [21] baseline. The results further highlight the importance of ORF identifiers, without which the performance even drops when switching to point level queries. This suggests that the provided element identifiers are essential for enabling effective point-level queries.
To investigate the viability of fusion approaches for the map decoder connection as proposed by TopoSD [28], we investigated a combination of direct cross-attention to the SD map encoder tokens and cross-attention to additional SD map BEV features, shown as “+ BEV Ft.” in Table˜3. The SD map BEV features are rasterized similarly to PMapNet [9], but contain processed SD map encoder tokens instead of grayscale images. Despite class-level differences, this combined connection showed no overall performance gain. Our results indicate that, when evaluated on a geographic split with the more diverse labels of M3TR [8], regular cross-attention alone is sufficiently expressive on its own.
Table 2: Comparison of SDTagNet with recent state-of-the-art non SD map prior methods on the more limited label setting of [3], evaluated on the Argoverse 2 original split and the geographic split of [30]. * = only total mAP reported in the original paper. †: values taken from MapTracker [3].
Dataset: Argoverse 2 Near Range (60 m 30 m)
Method Split Epochs $\textbf{AP}_{\textbf{div}}$ $\textbf{AP}_{\textbf{bou}}$ $\textbf{AP}_{\textbf{ped}}$ mAP
HIMap [33] Og. Split 24 72.4 73.2 72.4 72.7
MapUnveiler [11] Og. Split 30 74.2 71.9 72.5 72.9
StreamMapNet [30] Og. Split 72 74.2 66.1 70.5 70.3
MapTracker [3] Og. Split 35 80.0 73.7 77.0 76.9
SDTagNet Og. Split 24 81.7 76.3 76.1 78.0
StreamMapNet [30] [30] Geo Split 72 68.2 63.2 61.8 64.4
MapTracker [3] [30] Geo Split 35 75.1 68.9 70.0 71.3
SDTagNet [30] Geo Split 24 72.0 67.5 64.0 67.8
Far Range (100 m 50 m) from [30]
MapUnveiler [11] Og. Split 30 67.9 62.6 71.7 67.4
StreamMapNet [30] Og. Split 30 -* -* -* 57.7
SDTagNet Og. Split 24 80.2 72.9 81.7 78.3
StreamMapNet [30] [30] Geo Split 72 56.1 47.5 60.1 54.6
MapTracker [3] [30] Geo Split 35 64.6 58.5 71.2 64.8
SDTagNet [30] Geo Split 24 69.5 68.0 75.1 70.9
Base Enc.: SMERF [21] (all info.) AP
Pt. Lvl. NLP Emb. ORF Id. + BEV Ft. mAP dsh. sol. bou. cen. ped.
- - - - 45.9 38.0 54.4 50.0 48.5 38.2
- - - 44.3 34.9 53.0 49.5 45.5 38.7
- - - 45.8 37.0 55.4 47.9 49.2 39.3
- - 46.3 39.3 53.3 49.6 48.5 40.8
- - 45.0 33.6 53.2 50.2 47.7 40.6
- 48.1 36.0 55.2 53.3 52.6 43.3
48.1 38.2 56.3 52.6 50.7 42.3
@@
Table˜4 shows the evaluation results on the nuScenes dataset. The relative improvement of SDTagNet in the far range setting is even stronger, increasing performance compared to MapTRv2 by +4.1 mAP (+105%) and compared to existing methods by +2.1 mAP (+35%). This coincides with the general drop in mAP for all methods on the nuScenes geo split [18] and shows the benefits of our approach even on smaller datasets with less generalization capability. SMERF outperforms PMapNet here, indicating that these methods may be more susceptible to dataset differences than SDTagNet, which shows similar performance benefits in both cases. In the near range, the results mostly mirror those on Argoverse 2, with no method able to provide a significant performance gain. On nuScenes in particular, we observed some significant differences between the roads and structures contained in the SD map and the ground truth HD map (see Section˜5). We hypothesize this is one of the main causes of the lacking performance in the near range, where the visibility of road features is much higher and the model is less reliant on SD map features.
@@
Dataset: nuScenes Near Range (60 m 30 m)
@@ 5 Limitations
Despite SDTagNet’s strong performance, we observe several limitations. First, like all SD map prior methods, SDTagNet assumes congruency of SD map and HD map ground truth. Depending on the level of detail and quality of dataset ground truth and SD map, a number of map discrepancies can be observed. This is especially the case for crowdsourced databases like OSM, as shown for Argoverse 2 by [20] and nuScenes by [9]. We also noted significant differences on nuScenes, particularly concerning structures and roads that are missing from the ground truth HD map. Figure˜6(b) shows examples of a tunnel and smaller service roads that are not included in the nuScenes map, impacting model accuracy and training data quality.
Another limitation concerns the case of one-way roads. SDTagNet is able to correctly identify one-way roads from the SD map tag information, however in some cases the driving direction is predicted wrongly. An example of this is displayed in Figure˜6(a). Since OSM encodes one-way direction via polyline point order, this likely stems from the lack of point order information in the element identifiers. Adding point order identifiers for polyline features could help remedy this limitation.
@@
Figure 6: Examples for observed limitations of SDTagNet in Argoverse 2 (left) and nuScenes (right). Figure˜6(a): Even when correctly recognizing a one-way road, SDTagNet can fail to predict its correct driving direction. Figure˜6(b): Discrepancies between dataset HD maps and OSM SD maps like missing tunnels (left) or small service roads (left, right).
@@ 6 Conclusion
7 Acknowledgements
The authors gratefully acknowledge the computing time provided on the high-performance computer HoreKa by the National High-Performance Computing Center at KIT (NHR@KIT). This center is jointly supported by the Federal Ministry of Education and Research and the Ministry of Science, Research and the Arts of Baden-Württemberg, as part of the National High-Performance Computing (NHR) joint funding program (https://www.nhr-verein.de/en/our-partners). HoreKa is partly funded by the German Research Foundation (DFG). SD map data copyrighted OpenStreetMap contributors and available from https://www.openstreetmap.org.
@@ Appendix A Technical Appendices and Supplementary Material
The supplementary material focuses on three key aspects: Additional ablation studies on all individual components of SDTagNet (section˜A.1) and additional experiments on a non-geo split (section˜A.4), implementation details of the baselines and NLP pretraining (section˜A.2 and section˜A.3) as well as more qualitative examples and visualizations (section˜A.7).